OpenStudy (joshoyen):

Find the absolute value of each complex number. 1. |8-8i| 2. |3-6i| 3. |-4+3i|

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle |a+bi|=\sqrt{a^2+b^2} }\)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle |8+8i|=\sqrt{8^2+8^2} =\sqrt{2\times 8^2}=\sqrt{2}\times \sqrt{8^2}=? }\)

OpenStudy (solomonzelman):

Note (-8)^2 is same as 8^2.

OpenStudy (solomonzelman):

igtg, gb, good luck@!

OpenStudy (joshoyen):

Thanks!